I am stuck on a problem. My algebra teacher told us that we have to check it and show all our work and steps, and I can't get it right.
2(x-5+3x)=14
I've tried...
2(x-5+3x)=14
2x-10+6x=14
2x-2x-10+6x-2x=14
-10+4x=14
-10+10+4x=14+10
4x=24
4x/4=24/4
x=6
Then, when I check it I must be doing it wrong or don't have the right answer.
2(6-5+3*6)=14
12-10+6*12=14
12-10+72=14
2+72=14
74=14
It's supposed to be 14 on both sides, but I do not understand it. All of my other problems I got right, but I just for some reason do not understand this question. Please help me as soon as possible, thanks.

Answer by Alan3354(69443)

2(x-5+3x)=14
I've tried...
2(x-5+3x)=14
2x-10+6x=14
2x-2x-10+6x-2x=14
----------
You subtracted 2x from the left side, but not the right side.
----------
2(x-5+3x)=14
Divide by 2
(x-5+3x) = 7
4x-5 = 7
Add 5
4x = 12
x = 3